### v3

# Random Exclusion Codes: Quantum Advantages of Single-Shot Communication

Abstract
Useful applications of quantum information technologies can be found by identifying tasks in which quantum resources outperform their classical counterparts. In this work, we introduce a two-party communication primitive, random exclusion code (REC), which is a single-shot prepare-and-measure protocol where a sender encodes a random message into a shorter sequence and a receiver attempts to exclude a randomly chosen letter in the original message. We present quantum advantages in RECs in two ways: probability and dimension. We show that RECs with quantum resources achieve higher success probabilities than classical strategies. We verify that the quantum resources required to describe detection events of RECs have a smaller dimension than classical ones. We also show that a guessing counterpart, random access codes (RACs), may not have a dimension advantage over classical resources. Our results elucidate various possibilities of achieving quantum advantages in two-party communication.
I Introduction
One of the key challenges in quantum information theory is to address information tasks that demonstrate the distinct nature of quantum and classical resources. A gap in performance appearing from the distinction, even if it is small, may be amplified toward practical applications, namely, quantum advantages in communication winter. It is hence of both fundamental and practical importance to identify a simplest information scenario, such as a point-to-point communication protocol, such that quantum advantages are present. The Holevo theorem, concerning the channel capacity with i.i.d. quantum states, was the first attempt along these lines but failed to show a quantum advantage Holevo2011; 651037; PhysRevA.56.131.
It is random access codes (RACs) that have verified the existence of quantum advantages in two-party communication Ambainis2009; Ambainis2024RAC; Tavakoli2015; carmeli2020quantum. We emphasize that the setting is a single-shot scenario, in which a single outcome is obtained from a detection event, from which a receiver learns about a message. The guess is then correct with some probability. The quantum advantage is demonstrated by the fact that the guessing probability in RACs is higher if quantum, rather than classical, resources are used.
Although the connection is not direct, RACs are closely related to quantum state discrimination Helstrom:1969aa; Bae_2015; Barnett:09; Bergou:2007aa. Interestingly, quantum state exclusion, also known as anti-discrimination, which efficiently excludes a state from an ensemble, has been used as an effective tool to clarify the distinctions between quantum and classical resources in single-shot experiments Caves2002; PuseyPBR2012; PhysRevA.89.022336; HeKe18. Quantum advantages can be obtained in terms of a success probability PhysRevLett.125.110402; PhysRevResearch.2.033374, and a quantum-classical gap is also present in detection events. The Pusey-Barrett-Rudolph (PBR) theorem PuseyPBR2012, for instance, rules out a certain interpretation of quantum states by pinpointing a detection event that cannot happen with quantum resources.
Upon examining the fundamental tasks closely, it becomes clear that discrimination and exclusion are distinct in nature. Perfect discrimination of states may achieve perfect exclusion by reassigning events to different labels. The converse does not generally hold. Quantum state exclusion is thus operationally more general. State exclusion may therefore lead to greater possibilities of quantum advantages.
As discussed above, efforts to distinguish quantum and classical resources in a single-shot scenario take one of several approaches. One is to compare the success probability for a single detection event. As with RACs, the probability with quantum resources may be higher than that with classical resources. The other approach is to take into account detection events. Along these lines, the PBR theorem signifies a detection event that is impossible with quantum resources but is possible with independent ontic preparations PuseyPBR2012; Harrigan:2010aa. Note that the Bell theorem is comparable to the latter in that it also rules out an ontic interpretation of quantum states; in contrast to the PBR result, two parties repeat a Bell experiment for estimating joint probabilities PhysicsPhysiqueFizika.1.195; Brunner2014. It is worth mentioning that the quantum-classical gap from the Bell theorem has been developed as a practical tool for device-independent quantum information processing, such as randomness extraction Pironio:2010aa, quantum key distribution PhysRevLett.98.230501 and state certification Supic2020selftestingof; Supic2021deviceindependent. One would therefore expect useful applications to follow from other quantum-classical distinctions.
In this work, we present random exclusion codes (RECs) that leverage quantum state anti-discrimination to a communication primitive. Throughout, an $(n,m)$ REC of interest is a single-shot and prepare-and-measure protocol where a sender, Alice, encodes a random message of length $n$ using a set of $m$ letters into a $d$-dimensional system and transmits it. Bob, a receiver, realizes a decoding to exclude a chosen letter in the original message.
We show the quantum advantages of RECs in two ways: one is the probabilistic advantage for a single detection event, and the other is the dimension advantage for a collection of possible detection events. Firstly, a higher success probability in the exclusion task of a REC can be achieved with quantum resources than with classical ones. Secondly, quantum resources in a smaller Hilbert space achieve RECs that can be realized only in a larger classical dimension. Note that a classical dimension denotes the minimal number of distinguishable letters of an alphabet.
The importance of the results is twofold. On the one hand, we relate state exclusion, a fundamental tool for verifying distinctions between the elements of quantum and classical theories Caves2002; Bandyopadyhyay2014Exclusion; Crickmore2020Elimination; PuseyPBR2012, to a communication primitive that may find practical applications. On the other hand, we demonstrate the quantum-classical gap within a single-shot framework by highlighting the differences between quantum and classical supports for detection events, see also PuseyPBR2012. We also find that RACs may not have a comparable dimension advantage.
The article is structured as follows. We first remind the reader of some relevant results in the theory of quantum state exclusion and random access coding. We then introduce random exclusion codes and show that quantum resources outperform classical ones in the simplest non-trivial setting. We then analyse the required dimensional supports for perfect quantum and classical RECs and see that the former require fewer dimensions. Throughout, our results are contrasted with those from the equivalent RAC.
II Background
II.1 Quantum state exclusion
In information processes, one is typically interested in determing that a datum takes a particular value from a known, discrete alphabet. Exclusion conversely refers to the opposite task of determining that a datum does not have a given value Crickmore2020Elimination. The latter task is in general easier than the former, but recent works have shown that perfect implementation of this task is a signature of quantum behaviour.
In quantum state exclusion, a classical alphabet of size $m$ will be associated with a set of quantum states $\{\rho_{\alpha}\}_{\alpha=0}^{m-1}$ by a party Alice. She is given a value $\alpha$ at random and sends a system prepared in the appropriate state to a receiving party, Bob, who must announce one of the values which $\alpha$ is not. He determines his choice by implementing an $m$-outcome probability operator valued measurement (POVM) $\{\pi_{\beta}\}_{\beta=0}^{m-1}$.
We can distinguish between different criteria for state exclusion. Perfect state exclusion is that with zero-error, so that, whenever Alice prepares the state $\alpha$, Bob is guaranteed to announce a value $\beta\neq\alpha$. Uniform state discrimination, on the other hand, refers to the case in which all allowed outcomes occur with uniform probability. We will return to this point in later sections of the manuscript.
Figure 1: Trine states (black) in Eq. (1) and an anti-trine measurement (blue) in Eq. (2) can realize an unbiased exclusion task. For instance, an outcome $\pi_{0}$ concludes state $\psi_{1}$ or state $\psi_{2}$ with equal probabilities $1/2$ while excluding $\psi_{0}$ unambiguously.
The simplest non-trivial construction which is able to implement perfect and uniform state exclusion is the set of trine states
$|\psi_{\alpha}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle+e^{i2\pi\alpha/3}|1\rangle\right)\,\,\,\alpha=0,1,2\,$, (1)
and anti-trine measurements
$\pi_{\beta}=\frac{2}{3}|{\psi}_{\beta}^{\perp}\rangle\langle{\psi}_{\beta}^{\perp}|,\,\,\,{\beta}=0,1,2$ (2)
where $\langle\psi_{\beta}|{\psi}_{\beta}^{\perp}\rangle=0$. This set of states and measurements is shown in Fig. 1 and has been used to implement key distribution Phoenix2000 and in showing gaps between classical and quantum behaviour Heinosaari2024Simple. The defining feature in this example is symmetry. In similar symmetric constellations one finds the excluding POVM as a group covariant POVM HeKe18.
The problem of state exclusion was first studied by Caves et al. Caves2002 and came to wider attention when it was used to prove the PBR theorem regarding the $\psi$-onticity of quantum theory. Subsequent investigations have aimed to characterise the set of quantum ensembles which states can be unambiguously excluded Johnston2025tightbounds, and related conclusive exclusion to various foundational problems in quantum theory Havlicek2020Exclusion; Srikumar2024contextuality; Stratton2024Choi and the construction of zero-error communication channels Duan2016; Chiribella2025. The key takeaway from these results has been that perfect state exclusion is a signature of quantum behaviour and, furthermore, may be exploited for improved communications.
II.2 Random Access Codes
Random access codes are a communication primitive widely studied in the context of quantum theory for both their information-theoretic applications and foundational significance Ambainis2009; Ambainis2024RAC; Tavakoli2015; carmeli2020quantum. Alice holds a string $a=a_{1}...a_{n}$ consisting of $n$ independent classical labels each taken equiprobably from an $m$-symbol alphabet $\{0,\cdots,m-1\}$. This data is encoded into a system, either classical or quantum, of dimension $d$ which is then sent to the receiver, Bob. Bob’s task is to read out some subset of the information. He is given a question label, $i$, and aims to respond with $b_{i}$ such that $b_{i}=a_{i}$. Measures of his ability to perform this task use either the worst-case probability of success or the average success probability, given by
${p}^{\mathrm{(RAC)}}=\frac{1}{nm^{n}}\sum_{a}\sum_{i=1}^{n}\leavevmode\nobreak\ \mathrm{Prob}[b_{i}=a_{i}].$ (3)
If the communication system is quantum, Alice’s encoding will map the data to a set of density matrices $\{\rho_{a}\}$ and Bob’s decoding will be a set of positive operator-valued measures (POVMs) $M_{i}$. If the system is instead classical, then the encoding and decoding are simply maps to and from dits. In this article, we are particularly interested in the case in which $n=2$ and the quantum and classical systems to be compared are both two-dimensional. In this restricted setting, upper bounds are known to be Tavakoli2015; Ambainis2024RAC:
$\begin{split}\max_{Q}p^{(RAC)}&=\frac{1}{2}(1+\frac{1}{\sqrt{m}})\\ \max_{C}p^{(RAC)}&=\frac{1}{2}(1+\frac{1}{m}).\end{split}$ (4)
where the former optimisation is over quantum strategies and the latter over classical ones. In this case, and in general, it holds that $\max_{Q}p^{(RAC)}>\max_{C}p^{(RAC)}$.
III Random Exclusion Codes
We begin with the notation for an $(n,m)$ REC. A message of Alice’s is written by $a=a_{1}...a_{n}$ where each is chosen from a set of $m$ distinguishable letters, $a_{i}\in\{0,\cdots,m-1\}$. An encoding prepares a $d$-dimensional system, and we let $b_{j}$ denote Bob’s decoding of $a_{j}$ once he decides to read out $a_{j}$, see also Fig. 2. With an REC, Bob wants to find $b_{j}$ that is not equal to $a_{j}$. In this case, the probability of successfully excluding a correct value on average is given,
${p}^{\mathrm{(REC)}}=\frac{1}{nm^{n}}\sum_{a}\sum_{i=1}^{n}\leavevmode\nobreak\ \mathrm{Prob}[b_{i}\neq a_{i}].$ (5)
A quantum realization signifies a preparation of a $d$-dimensional quantum state and $m$-outcome measurements for Bob to obtain $b_{j}$.
For RACs, in contrast to RECs, Bob wants to guess Alice’s message and the average success probability is computed from Eq. (3). In both the tasks, above in Eqs. (3) and (5), parties aim to maximize the success probability over all strategies with given resources.
Note that state discrimination and state exclusion are equivalent for an alphabet of two letters: guessing one immediately excludes the other, and vice versa. For example, for binary messages $a_{i}\in\{0,1\}$, Bob can apply two-state discrimination and it holds that
$\max_{S}{p}^{\mathrm{(RAC)}}=\max_{S}{p}^{\mathrm{(REC)}},$ (6)
where the maximization runs over all strategies $S$. From this, one immediately concludes that RECs bear quantum advantages of RACs in the $(2,2)$ instance. However, the relation above no longer holds when an alphabet contains more than two letters as excluding a correct value does not yield a single outcome but rather multiple ones.
Let us then consider the next simplest case, the $(2,3)$ REC, where two-trit messages $a=a_{1}a_{2}$ for $a_{i}\in\{0,1,2\}$ are encoded to a binary system, for which a quantum (classical) realization is a single qubit (bit). In what follows, we demonstrate a quantum advantage by showing that Bob has a higher probability of excluding a message sent by Alice with quantum rather than classical resources.
IV Quantum advantage in $(2,3)$-RECs
We prove analytically that, if the communication system is a classical bit, the maximum value of the average probability of success in the considered task is $\frac{8}{9}$. We then show that if instead quantum systems of dimension two are used, a larger value of $\frac{7+\sqrt{2}}{9}$ can be attained. It is worth mentioning here that for the considered REC, since the alphabet $\{0,1,2\}$ is non-binary, the corresponding RAC task is not equivalent, which can be seen from the fact that the maximum classical value for the RAC task is $\frac{5}{9}$ (see Result 3 in the Ref. deba+PRA2023). The inequivalence also holds for the quantum values, where the maximum success probability for the corresponding RAC is $\frac{4+\sqrt{2}}{9}$.
IV.1 Optimal bit implementation
We first compute the maximum average success probability with classical resources, which can be summarised as follows:
Proposition 1.
The maximum classical average success probability in the $(2,3)$ random exclusion task, consisting of length $n=2$ words using a size $m=3$ alphabet and $d=2$ dimension communicated system, is $8/9$.
Proof.
Let us begin by calculating the optimal value over all deterministic protocols. The two stages of a deterministic protocol are as follows: (i) The set of all possible words $\{a_{1}a_{2}:a_{1},a_{2}\in\{0,1,2\}\}$ is partitioned into two parts $\{\mathbb{P}_{0},\mathbb{P}_{1}\}$; if a word to Alice belongs to the part $\mathbb{P}_{0}$ ($\mathbb{P}_{1}$), she communicates to Bob $c=0$ ($c=1$). (ii) Bob, on receiving a question $i\in\{1,2\}$, answers $b_{i}^{c}\in\{0,1,2\}$ following a deterministic strategy, i.e., a map $\{b_{1}^{0},b_{2}^{0},b_{1}^{1},b_{2}^{1}\}\mapsto\{0,1,2\}$.
Let us first note that, for a given deterministic strategy of Bob, two specific words follow, namely, the word $b_{1}^{0}b_{2}^{0}$ and the word $b_{1}^{1}b_{2}^{1}$. Now, let us consider that Alice receives either of these words. The message she communicates will then be determined by her partition, and there are four cases to consider:
  1. 1.
    $b_{1}^{0}b_{2}^{0},b_{1}^{1}b_{2}^{1}\in\mathbb{P}_{0}$,
  2. 2.
    $b_{1}^{0}b_{2}^{0},b_{1}^{1}b_{2}^{1}\in\mathbb{P}_{1}$,
  3. 3.
    $b_{1}^{0}b_{2}^{0}\in\mathbb{P}_{0}$ and $b_{1}^{1}b_{2}^{1}\in\mathbb{P}_{1}$,
  4. 4.
    $b_{1}^{0}b_{2}^{0}\in\mathbb{P}_{1}$ and $b_{1}^{1}b_{2}^{1}\in\mathbb{P}_{0}$.
We study each case in turn. In Case 1, when Alice receives the word $a=b_{1}^{0}b_{2}^{0}$ she will communicate the value $c=0$ to Bob. For each question $i$ that he is asked he will output $b_{i}^{0}$, which was the value he aimed to exclude. Therefore, the protocol fails two times. Note that, in this case, if instead she received the message $a=b_{1}^{1}b_{2}^{1}$, Bob will end up answering $b_{1}^{0}$ and $b_{2}^{0}$ which need not be the same as the initial string; he can therefore succeed in these cases. Case 2 follows the same argument as Case 1. For Case 3, when Alice receives $a=b_{1}^{0}b_{2}^{0}$, her message is $c=0$ and Bob’s output will therefore be $b_{i}^{0}$, so he fails for both questions. Similarly, he will be incorrect whenever $b_{1}^{1}b_{2}^{1}$ is the message received by Alice. Thus, in Case 3, Bob fails in at least four instances of the protocol.
In Case 4, as in Case 3, it can never follow that $b_{1}^{0}b_{2}^{0}=b_{1}^{1}b_{2}^{1}$ as this would mean that Alice puts the same word into two sides of the partition. Let us first suppose that $b_{1}^{0}b_{2}^{1}\neq b_{2}^{0}b_{1}^{1}$. Then, when Alice receives $a=b_{1}^{0}b_{2}^{1}$ she will communicate either $c=0$ or $c=1$ depending on which partition it belongs to. In the former case, Bob’s answer is in the string when his question is $i=1$ and in the latter case it is in the string when his question is $i=0$. Similarly, he will be incorrect in at least one position $i$ when the initial string is $a=b_{2}^{0}b_{1}^{1}$. Case 4 therefore fails at least four times if $b_{1}^{0}b_{2}^{0}=b_{1}^{1}b_{2}^{1}$. The final situation is that $b_{1}^{0}b_{2}^{1}=b_{2}^{0}b_{1}^{1}$. This implies further that $b_{1}^{0}=b_{2}^{0}$ and $b_{1}^{1}=b_{2}^{1}$. On considering the words $a=b_{1}^{0}b_{1}^{1}$ and $a=b_{1}^{1}b_{1}^{0}$ , irrespective of which partition they belong to the protocol will fail at least once for each. Therefore, in Case 4 as well it is observed that the protocol fails to exclude in at least two instances.
We can hence see that, in all four cases, for any deterministic strategy Bob answers incorrectly in at least two out of eighteen instances of the protocol. The maximal average success probability with deterministic strategies is therefore upper bounded by ${p}^{\mathrm{(REC)}}\leq 16/18=8/9$. This bound is, in fact, tight and can be achieved if Alice uses the partition $\mathbb{P}_{0}=\{11,12,21,22\}$ and $\mathbb{P}_{1}=\{00,01,02,10,20\}$, and if Bob’s decoding is $b_{i}^{c}=c$. One can easily verify that this strategy gives the desired value.
Finally, we show that probabilistic strategies cannot perform better. Let’s assume that there are $K$ deterministic strategies $S_{j}$ where $j\in\{1,\cdots,K\}$ and the parties have access to shared random variables $\{\lambda_{j}\}_{j=1}^{K}$ with probability distribution $p(\lambda_{j})=p_{j}$. Note that any probabilistic strategy can be implemented with access to these random variables. If Alice and Bob apply strategy $S_{j}$ with probability $p_{j}$ we get a probabilistic strategy $S=\{p_{j},S_{j}\}_{j=1}^{K}$. The average success probability with any such strategy $S$ is ${p}^{\mathrm{(REC)}}_{S}=\sum_{j}p_{j}{p}^{\mathrm{(REC)}}_{S_{j}}\leq\sum_{j}p_{j}(8/9)=8/9$. This completes the proof. ∎
Figure 2: In an $(n,m)$ REC, a sender prepares a message, $a=a_{1}\cdots a_{n}$, of length $n$ where each is chosen from an alphabet of $m$ letters $a_{i}\in\{0,1,\cdots,m-1\}$. An encoding corresponds to a preparation of a single system of cadinality $d$, which is sent to Bob, who first decides which letter he will read, say the $j$-th one, and reads $b_{j}$. Bob wants to exclude the message sent by Alice, i.e., $b_{j}\neq a_{j}$. In RACs, parties want to have $b_{j}=a_{j}$. For these purposes, parties optimize strategies with quantum or classical resources.
IV.2 Qubit implementation and quantum advantage
We next present a $(2,3)$ REC with single qubit states in the following,
$\rho_{a}:=\rho_{a_{1}a_{2}}=|\psi_{a_{1}a_{2}}\rangle\langle\psi_{a_{1}a_{2}}|\leavevmode\nobreak\ \mathrm{for}\leavevmode\nobreak\ a_{i}\in\{0,1,2\}$ (7)
where
$|\psi_{00}\rangle$ $=$ $\frac{|0\rangle+e^{i\pi/4}|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ |\psi_{01}\rangle=\frac{|0\rangle+e^{-i\pi/4}|1\rangle}{\sqrt{2}},$
$|\psi_{10}\rangle$ $=$ $\frac{|0\rangle+e^{i3\pi/4}|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ |\psi_{11}\rangle=\frac{|0\rangle+e^{-i3\pi/4}|1\rangle}{\sqrt{2}},$
$|\psi_{20}\rangle$ $=$ $\frac{|0\rangle+e^{i\pi/2}|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ |\psi_{21}\rangle=\frac{|0\rangle+e^{-i\pi/2}|1\rangle}{\sqrt{2}}$
$|\psi_{02}\rangle$ $=$ $\frac{|0\rangle+|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ |\psi_{12}\rangle=\frac{|0\rangle-|1\rangle}{\sqrt{2}},\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ \rho_{22}=\mathbb{I}/2.$
We construct a decoding with two-outcome measurements,
$M_{1}=\{M_{10},M_{11},M_{12}\}\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ M_{2}=\{M_{20},M_{21},M_{22}\}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\$(8)
where
$M_{10}=|+\rangle\langle+|,\leavevmode\nobreak\ M_{11}=|-\rangle\langle-|,\leavevmode\nobreak\ M_{12}=0$
$M_{20}=|+i\rangle\langle+i|,\leavevmode\nobreak\ M_{21}=|-i\rangle\langle-i|,\leavevmode\nobreak\ \mathrm{and}\leavevmode\nobreak\ M_{22}=0,$
and $|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$ and $|\pm i\rangle=(|0\rangle\pm i|1\rangle)/\sqrt{2}$. Notice that the states and measurements above are on a great circle of the Bloch sphere. We display the states and measurements of our scheme in in Fig. 3.
Figure 3: Illustration of states and measurements for the scheme discussed in text. States are displayed in light blue and the measurements $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ in red and green respectively.
The quantum protocol works as follows. Alice sends a state $\rho_{a_{1}a_{2}}$ and Bob first decides which one, $a_{1}$ or $a_{2}$, he is going to read. For the first trit $a_{1}$, he applies measurement $M_{1}$, and for the second, $a_{2}$, he applies $M_{2}$. The probability of successfully excluding a one-bit message of Alice is given by,
$p^{\mathrm{(REC)}}_{Q}$ $=$ $\frac{1}{18}\sum_{a}\sum_{i=1}^{2}{\rm Tr}[(M_{i(a_{i}\oplus 1)}+M_{i(a_{i}\oplus 2)})\leavevmode\nobreak\ \rho_{a_{1}a_{2}}]\leavevmode\nobreak\ \leavevmode\nobreak\$(9)
where $\oplus$ represents addition modulo three. It follows that
$p^{\mathrm{(REC)}}_{Q}=\frac{7+\sqrt{2}}{9}$ (10)
which is larger than the maximum classical value $8/9$ (Proposition 1.). Thus, the utility of quantum resources in RECs is demonstrated.
IV.3 Maximum quantum advantage
There may exist states and measurements other than those in Eqs. (7) and (8), in particular using POVMs containing no null elements, such that the success probability is even higher. That is, one can ask if the bound in Eq. (10) is maximal. Our answer is in the affirmative. Firstly, we find that to maximize the success probability with qubit transmission, it suffices to restrict states and measurements to a half-plane on a Bloch sphere. This also holds true even if parties aim at a guessing task, and hence, applies to RACs with qubits. Secondly, extensive numerical optimization of the average success probability finds Eq. (10) with two POVM elements are zeros. We detail the calculations in Appendix LABEL:appendA. The gap between quantum and classical bounds is
$\Delta_{QC}^{(\mathrm{REC})}=\max_{Q}p^{(\mathrm{REC})}-\max_{C}p^{(\mathrm{REC})}=\frac{\sqrt{2}-1}{9},$ (11)
which is approximately $0.046$.
For comparison, let us also consider a $(2,3)$ RAC with single-qubit transmission. An upper bound to success probabilities with classical strategies is known as $5/9$ deba+PRA2023. We compute the maximum average probability with quantum resources, applying the aforementioned result of reducing to a great circle. We find that a relabelling of our quantum protocol for the REC gives the quantum maximum for the RAC (see Appendix LABEL:appendA); messages $a=a_{1}a_{2}$ are encoded in the states orthogonal to ones given in Eq. (7) and the same measurements as in Eq. (8) are applied for decoding. The quantum bound is obtained as $(4+\sqrt{2})/9$. Interestingly, the gap is identical to Eq. (11), i.e.,
$\Delta_{QC}^{(\mathrm{RAC})}=\Delta_{QC}^{(\mathrm{REC})}$ (12)
for the considered $(2,3)$ RAC and $(2,3)$ REC.
To summarise, we have characterised bit and qubit implementations of the $(2,3)$ REC and shown that using quantum systems allows for a higher average probability of success. We also contrasted the behaviour of the $(2,3)$ RAC.
V Dimensional Advantages
In what follows, we are motivated by the observation that a single-shot scenario gives one outcome among possible detection events. This raises the question of which dimensional resources are required to observe distinctions between classical and quantum behaviours in the single-shot setting. Having characterised above the optimal performance of the $(2,3)$ REC in two-dimensional systems, we now turn our attention to a more general setting. We investigate the minimal quantum and classical resources to describe possible detection events in an errorless $(2,3)$ REC. We first note that there is no difference in behaviour for the standard scenario. This further motivates us to add an extra constraint, uniformity of outcomes, to our protocol. Classical implementations of this modified protocol require nine-dimensional systems, whereas quantum systems of four-dimensional behaviour, showing a dimensional advantage in uniform RECs. This behaviour is contrasted with RACs, where no such advantage can be observed.
V.1 Perfect random exclusion codes
We first note that no dimensional advantage can be observed for zero-error implementations of the basic $(2,3)$ REC which has been discussed so far. To do this, we note that there exists a three-dimensional classical, and therefore also quantum, strategy which may implement this task with no error. The explicit strategy is for Alice to encode her messages using the partition
$\begin{split}\mathbb{P}_{0}&=\{11,12,22\}\\ \mathbb{P}_{1}&=\{00,02,20,22\}\\ \mathbb{P}_{2}&=\{01,10\}\end{split}$ (13)
and send the relevant trit value to Bob. If he announces the value he has received, he will exclude without error. As we saw in the previous section, there exists neither a classical nor a quantum strategy that is able to achieve an average success probability of one using a two dimensional implementation. We can thus conclude that three is the minimal dimension requirement in both cases.
V.2 Uniform random exclusion codes
In order to witness a dimensional advantage in the $(2,3)$ REC, we must therefore analyse a modified form of the problem. This extra requirement will be uniformity of the outcomes, meaning that the correct answers by Bob all occur with equal probability. This constraint is regularly considered in the literature on exclusion tasks (see, e.g., Ref Heinosaari2024Simple; PhysRevResearch.2.033374) in order to avoid certain trivial solutions to the problem of state exclusion. It is in fact uniform state exclusion that can be considered a strong signature of quantum, rather than classical, behaviour. The constraint can be furthermore justified if we consider that Alice and Bob are trying to convince a referee, who provides Alice with her string, of their ability to perform quantum exclusion. If their results are non-uniform, then the referee will be able to use their strategy against them when the task is performed over several rounds. For example, in our optimal qubit scheme, Bob will never announce two as his answer. The referee can thus always send strings that do not contain this value, lowering their probability of success. To summarise, our aim here will be to determine the minimum quantum and classical dimensions required to implement a zero-error, uniform $(2,3)$ REC.
The natural framework for studying dimensional requirements is that of communication matrices Kerppo2023Communication, which we review here for the reader’s convenience. A communication matrix is a row-stochastic matrix encoding the conditional probabilities of prepare-and-measure experiments. If the inputs are a set of classical labels $\alpha\in[0,m-1]$ and the outputs a set of classical labels $\beta\in[0,n-1]$ then the communication matrix is an $m\times n$-sized row-stochastic matrix $C$ in which
$C_{\alpha\beta}={\rm Prob}_{M|P}(\beta|\alpha),$ (14)
where the subscripts $P$ and $M$ denote preparation and measurement respectively. For instance, the communication matrix of a binary state discrimination task in which two states are perfectly distinguished would be the diagonal matrix $S_{2}={\rm diag}[1,1]$. Similarly, a uniform state exclusion task with three states excluded in an unbiased manner would correspond to the communication matrix
$A_{3}=\frac{1}{2}\begin{bmatrix}0&1&1\\ 1&0&1\\ 1&1&0\end{bmatrix}$. (15)
This can be implemented in a quantum system by the trine-antitrine construction given earlier in Eqs. 1 and 2 .
The use of communication matrices in quantum information science regards the different sets $C^{d}_{cl}$ and $C^{d}_{q}$ which can be generated by $d$-dimensional classical and quantum systems respectively. We recall that the quantum dimension is that of the associated Hilbert space and the classical one is defined by the number of perfectly distinguishable letters in a communication channel. An important theorem due to Frenkel and Weiner Frenkel2015Classical tells us that the closure of the quantum and classical two sets is equal whenever they are generated by systems with matching dimension. The question becomes more intriguing if we ask how many dimensions are required of a quantum or classical system to implement a given communication matrix. With this in mind, it has been shown that the minimal quantum and classical dimensions follow from two different properties of a given communication matrix. Firstly, it is known that the minimum classical dimension required is given by the nonnegative rank:
Definition 1.
Nonnegative rank. A matrix $C$ is said to be nonnegative if all elements are nonnegative. The nonnegative rank Cohen1993Nonnegative of an $n\times m$ matrix $C$, denoted $\mathrm{rank}_{+}(C)$, is defined as the smallest number $k$ such that there exists an $n\times k$ nonnegative matrix $L$ and $k\times m$ nonnegative matrix $R$ such that $LR=C$.
Similarly, the minimum dimension of a quantum system is given by the positive semi-definite rank: